two real numbers that have a sum of 5 and the same two numbers that multipled equal 5.
Answer by mananth(16949)

x+y =5
xy = 5
x=5/y
substitute x in first equation
5/y +y =5
multiply equation by y
5+y^2=5y
y^2-5y+5=0
y^2-5y=-5
complete the square on the left side
y^2-5y+25/4=-5+25/4
(y-5/2)^2= 5/4
take the square root
y-5/2 = +/- sqrt(5)/2
y= 5/2 + sqrt(5)/2 = (5+sqrt5)/2
y=3.618
OR
y= 5/2 -sqrt(5)/2
y= (5-sqrt(5))/2
y=1.38